OpenStudy (anonymous):

The probability that an airplane engine will fail in a transcontinental flight is 0.002. Assuming that engine failures are independent of each other, what is the probability that on a certain transcontinental flight, a six-engine plane will experience: (a) Exactly one engine failure? (Round your answer to five decimal places.) Correct: Your answer is correct. (b) Exactly two engine failures? (Round your answer to six decimal places.) I NEED THIS ONE (c) More than two engine failures? (Note: In this event, the airplane will crash. Round your answer to nine decimal places.)

OpenStudy (anonymous):

I need B

OpenStudy (anonymous):

Ok, let's see. 0.002 ^ 2 = 0.000004

OpenStudy (anonymous):

Do you also need C?

OpenStudy (anonymous):

no I have C

OpenStudy (anonymous):

is B that easy?

OpenStudy (anonymous):

Your answer to B is wrong @D35TR0Y3R . The answer is (0.002)^2 * (0.998)^4 * 6C2 = (0.002)^2 * (0.998)^4 * 15 = .00006

OpenStudy (anonymous):

That 6C2 is combinations of 6 things taken 2 at a time..or 6! / (6-2)! 2!

OpenStudy (anonymous):

Are you sure you got A correct? .002 * .998^5 * 6 = .01188

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

yes